OpenStudy (anonymous):

What does 5 mean? There is no fraction bar. 1

OpenStudy (saifoo.khan):

might be a mis print, lol.

OpenStudy (anonymous):

You can see it on this page. It's the 2nd figure down. I don't think it's a typo : / http://www.regentsprep.org/Regents/math/algtrig/ATS7/BLesson.htm

OpenStudy (saifoo.khan):

oh this is a matrix.

OpenStudy (saifoo.khan):

matrices*

OpenStudy (anonymous):

I don't understand how to solve it

OpenStudy (saifoo.khan):

which fig?

OpenStudy (anonymous):

the 2nd one. i don't understand how the matrices apply to the equation

OpenStudy (saifoo.khan):

lol, same here.

OpenStudy (anonymous):

for instance \[\left(\begin{matrix}5 \\ 1\end{matrix}\right) \left( 1/4 \right) (3/4)^4\]

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

It's a binomial coefficient.\[\binom{n}{k} = \frac{n!}{k!(n-k)!}\]where \[n! = 1 \cdot 2 \cdot 3 \cdot 4 \dotsm n,\]is the factorial function, that is, multiplying every natural number between 1 and n (for example, 2! = 1 · 2 = 2, 6! = 1 · 2 · 3 · 4 · 5 · 6 = 720). This number represents the amount of ways you can choose k elements from a set of n elements.